Disregard the work shown it’s not correct, I need correct answer with proper explanation
Thank you
12. (6 points) If the reaction below initially consisted of just carbon and 1.50 atm of CO2 what are the equilibrium pressures of C02 and CO? Kp 14.1 2CO(g) 1.56 C (s) + CO2(g) 1 ?? 2 x 1 a yv>,14 , i v-?.is,0 Poo = 2a
Answer #1

                        C(s) + CO2(g) ------------>2CO

I                                    1.5                        0

C                                    -x                        2x

E                                  1.5-x                     2x

                    Kp    = P^2CO/PCO2

                    14.1   = (2x)^2/1.5-x

                    14.1(1.5-x) = 4x^2

                   21.15-14.1x -4x^2 = 0

                   4x^2+14.1x-21.15 = 0

                       x   = -14.1+\sqrt{}(14.1)^2 -4*4*-21.15/2*4

                       = -14.1 + \sqrt{} 537.21/8

                           = -14.1+23.17/8   = 1.134

PCO2   = 1.5-x   = 1.5-1.134   = 0.37atm

PCO    = 2x      = 2*1.134            = 2.27atm